OpenStudy (anonymous):

Let C be between D and E. Use the Segment Addition Postulate to solve for v. DC=5v-12 CE=6v-10 DE=22

OpenStudy (anonymous):

DE = DC + CE 22 = 5v-12 + 6v-10 11v = 22 + 22 = 44 => v = 44/11 = 4

OpenStudy (anonymous):

im confused on how you got 22 twice?

OpenStudy (anonymous):

Given DE = 22 ( you posted )

OpenStudy (anonymous):

no when you wrote 11v= 22 + 22=44.. how did you get the second 22?

OpenStudy (anonymous):

-10 -12 = -22 Calculation ... v= ...+ 22

OpenStudy (anonymous):

I dont understand if it equals to -22 how did it change to postive 22?